# Lagrangian chaos and unique ergodicity for stochastic primitive equations

### v1

@@ Abstract.
We show that the Lagrangian flow associated with the stochastic 3D primitive equations (PEs) with non-degenerate noise is chaotic, i.e., the corresponding top Lyapunov exponent is strictly positive almost surely. This result builds on the landmark work by Bedrossian, Blumenthal, and Punshon-Smith [16] on Lagrangian chaos in stochastic fluid mechanics. Our primary contribution is establishing an instance where Lagrangian chaos can be proven for a fluid flow with supercritical energy, a key characteristic of 3D fluid dynamics. For the 3D PEs, establishing the existence of the top Lyapunov exponent is already a challenging task. We address this difficulty by deriving new estimates for the invariant measures of the 3D PEs, which capture the anisotropic smoothing in the dynamics of the PEs. As a by-product of our results, we also obtain the first uniqueness result for invariant measures of stochastic PEs.
@@ 1. Introduction
In this paper, we consider the stochastic primitive equations (PEs) with additive noise on $\mathcal{O}=\mathbb{T}^{2}_{x,y}\times(0,1)$:
@@
Hence, the horizontal velocity field $v_{t}$ is periodic in all directions, while the Dirichlet condition on $w_{t}$ ensures the impermeability of the boundary $\partial\mathcal{O}$. In particular, $v_{t}$ and $w_{t}$ can be considered as maps on the three-dimensional torus $\mathbb{T}^{3}$.
The PEs (1.1) are one of the fundamental models for geophysical flows used to describe oceanic and atmospheric dynamics, where, in a typical region, the vertical scale is much smaller than the horizontal ones. In such domains, the PEs well-approximate the 3D Navier-Stokes equations. In the deterministic case (i.e., $Q=0$), rigorous justifications by means of the small aspect ratio limit can be found in [12, 32, 56]. Detailed information on the geophysical background for the various versions of the deterministic PEs can be found in, e.g., [52, 62, 69].
The PEs have been widely studied in recent years, therefore, it is impossible to give a complete overview here. However, we highlight some key references. In the deterministic setting, we should mention the seminal work by Lions, Temam, Wang [58, 59], as well as the breakthrough result by Cao and Titi [22] on global well-posedness in $H^{1}$. Extensions in various directions can be found in [34, 35, 43, 44, 49, 53]. In the stochastic setting, global well-posedness was first obtained by Debussche, Glatt-Holtz, Temam and Ziane in [27] (see also [26]). For further developments, the reader is referred to [2, 4, 5, 21, 28, 45] and the references therein.
In contrast to global well-posedness, the dynamic properties of solutions to PEs are far less understood. The primary goal of this manuscript is to prove that the Lagrangian trajectories associated with the solution $u_{t}$ of the stochastic PEs (1.1) are chaotic in the sense that they are highly sensitive to initial conditions. This phenomenon, often referred to as Lagrangian chaos, is established in the following.
@@
(1.3) $\frac{\mathrm{d}}{\mathrm{d}t}\phi^{t}(\mathbf{x})=u_{t}(\phi^{t}(\mathbf{x})),\qquad\phi^{0}(\mathbf{x})=\mathbf{x}\in\mathbb{T}^{3},$
@@
(1.4) $\lim_{t\to\infty}\frac{1}{t}\log|\nabla\phi^{t}(\mathbf{x})|=\lambda_{+}>0\text{ for all $\mathbf{x}\in\mathbb{T}^{3}$ with probability }1.$
The rigorous formulation of the above is given in Theorem 3.5 below. Equation (1.4) shows that nearby particles are exponentially separated by the flow $\phi^{t}$. The exponential expansion rate $\lambda_{+}$ is usually called the (top) Lyapunov exponent of $\phi^{t}$. Consequences on scalar turbulence of Theorem 1.1 or, more precisely, a version of (1.4) with $\nabla\phi^{t}(\mathbf{x})$ replaced by $(\nabla\phi^{t}(\mathbf{x}))^{-\top}$, are given in Corollary 3.7.
@@
A key element in the proof of Theorem 1.1 is the existence of a unique invariant measure for the stochastic PEs (1.1). Invariant measures play a crucial role in understanding the long-term behaviour of solutions to a given SPDE [25], and their uniqueness has significant implications, including ergodicity and the complete characterization of long-time dynamics through the invariant measure. For definitions, the reader is referred to Subsection 3.1. While the existence of ergodic measures for PEs was established by Glatt-Holtz, Kukavica, Vicol and Ziane in [37], their uniqueness remains an open problem (see the comments below Theorem C in [19] and the recent work [57]). The main obstacle in establishing uniqueness is the weak energy bounds available for PEs, cf., [37, Section II] and [5, Theorem 3.6 and Remark 3.10]. Here, the weakness is with respect to integrability in the probability space and invariant measures.
In this manuscript, by leveraging on new high-order energy estimates for (1.1) established in Section 5, we are able to overcome these challenges. The following is an informal statement and a special case of Proposition 3.2 and Theorem 3.3.
@@
To the best of our knowledge, Theorem 1.2 presents the first uniqueness result for invariant measures of PEs (1.1). Our proof of uniqueness follows the well-established strategy involving strong Feller and irreducibility properties of transition Markov kernels, as used in [31] for the 2D Navier-Stokes equations. For the relevant terminology, we refer the reader to Subsection 3.1 or [25]. In our setting, the high-order energy estimates established in Section 5 play a central role. It is worth noting that, due to the use of the strong Feller approach, we require that the noise acts on all Fourier modes, cf., (3.3) and Subsection 2.2.2. Establishing the uniqueness of invariant measures in the case of degenerate noise–where some Fourier modes are not randomly forced–appears to be extremely challenging due to the above-mentioned weak energy bounds. In particular, known methods, such as asymptotic couplings [38, 42] (see [38, Subsection 3.2] for the two-dimensional case of (1.1)), asymptotic strong Feller [40], or the $e$-property [51], do not seem to be applicable in this context.
@@ 1.1. Background and contribution
Recall that energy supercriticality is typical of three-dimensional models, and this produces weak energy bounds with respect to the probability space. Let us emphasize that the latter also appears in two-dimensional models with boundary noise (see Subsection 1.2 for more details).
The stochastic PEs (1.1) serve as a prototypical example of globally well-posed 3D fluid models in which both of the aforementioned conditions are simultaneously satisfied. While the boundary conditions are explicitly given in (1.2) (though the domain and conditions are relatively simple), the energy supercriticality can be seen from a scaling analysis; see [2, Subsection 1.2] for details. It is important to note that the PEs (1.1) possess the scaling of the 3D Navier-Stokes equations, and thus, they are energy supercritical.
The essential ingredients in the application of the general framework in [16, Section 4] to (1.1) can be (roughly) summarized as follows:
@@
    An integrability condition on ergodic measures of the fluid model.
@@
    Suitable approximate controllability results for:
    • The Jacobian process $(u_{t},\phi^{t}(\mathbf{x}),\nabla\phi^{t}(\mathbf{x}))_{t}.$
    • The projective process $(u_{t},\phi^{t}(\mathbf{x}),\nabla\phi^{t}(\mathbf{x})\xi/|\nabla\phi^{t}(\mathbf{x})\xi|)_{t}.$
In item (3), $\mathbf{x}\in\mathbb{T}^{3}$ and $\xi\in{\mathbb{R}}^{3}\setminus\{0\}$.
@@
(1.7) $w_{t}(\mathbf{x})=-\int_{0}^{z}\nabla_{x,y}\cdot v_{t}(x,y,z^{\prime})\,\mathrm{d}z^{\prime},\ \quad\mathbf{x}=(x,y,z)\in\mathbb{T}^{3}.$
Here, $\nabla_{x,y}\cdot$ is the divergence operator with respect to the horizontal variables $(x,y)$; see below for the notation. Consequently, (1.6) requires the integrability of quantities involving second-order derivatives of the horizontal velocity $v_{t}$. It is important to note that, in the existing literature, only logarithmic bounds on second-order derivatives are available; see [37, Theorem 1.9]. However, the latter is too weak to check (1.6), therefore ensuring the existence of the Lyapunov exponent (1.5).
In this work, we address these issues by using $L^{p}$-estimates and leveraging the anisotropic smoothing effect associated with the nonlinearity $w\partial_{z}v$, with $w$ as in (1.7). For further details, the reader is referred to Theorem 4.2 and the comments below it. Additionally, we obtain (1.6) with spatial integrability $L^{1}$ replaced by $L^{p}$ for any $p<\frac{4}{3}$. We also expect that the $L^{2}$-case of (1.6) does not hold.
@@
Before discussing the subsequent item, let us comment on other possible choices of smoothing for the Markov kernel of the Lagrangian process. In the recent work [23], the authors proved that in the context of the 2D Navier-Stokes equations the results in [16] (as well as subsequent works [13, 14, 15]) can be extended to the case where the noise acts on only finitely many Fourier modes (i.e., degenerate noise). This extension is achieved through the use of the asymptotic strong Feller property, which was introduced in [40] to study unique ergodicity of the 2D Navier-Stokes equations with degenerate noise. However, as noted below Theorem 1.2, it remains unclear how to establish asymptotic strong Feller for the stochastic PEs (1.1) itself in the case of degenerate noise. Therefore, the extension presented in [23] of the results in [16] to the case of degenerate noise does not currently seem applicable in the context of the stochastic PEs (1.1).
@@
To conclude, let us emphasize that, as in [16], we will obtain a stronger result than (1.4) in Theorem 1.1. Specifically, using the random multiplicative ergodic theorem [50, Theorem III.1.2] (also used in [16], see Proposition 3.16 therein), one can derive from Theorem 1.1 the following stronger result: For all $\xi\in{\mathbb{R}}^{3}\setminus\{0\}$,
(1.8) $\lim_{t\to\infty}\frac{1}{t}\log|\nabla\phi^{t}(\mathbf{x})\xi|=\lambda_{+}>0\text{ for all $\mathbf{x}\in\mathbb{T}^{3}$ with probability }1.$
@@
    Uniqueness of the invariant measure for the projective process.
The above item will be established through a weak irreducibility argument and the strong Feller property for the projective process, as described in Propositions 3.11 and 3.12. Let us emphasise that the strong Feller property for the projective process is a stronger result than what is required in (2) or the previous list.
@@ 1.2. Future challenges
  • Mixing and enhanced dissipation for passive scalars. As shown in the works [13, 14], the positivity of the top Lyapunov exponent of the flow map associated with a stochastic fluid model often implies mixing and enhanced dissipation results for passive and diffusive scalars, respectively. It is important to note that these results are stronger than the exponential growth for passive scalars presented in Corollary 3.7. While the positivity of the top Lyapunov exponent for the induced flow map is a key element in [13, 14], these results also rely on the existence of Lyapunov functions that provide quantitative control over high deviations of the velocity $u_{t}$, as discussed in [14, Subsection 2.4]. The construction of such Lyapunov functions remains an open problem for stochastic PEs (1.1).
  • Other geophysical flows: Boundary noise. The interaction between the wind and ocean surface plays a crucial role in ocean dynamics, see [62, 63, 36]. In the case of highly oscillating wind, and under the assumption of a time-scale separation, the wind’s effect can be idealized as noise acting on the boundary. Recent well-posedness results for fluid dynamics models with boundary noise are presented in [3, 6, 18]. The physical relevance of extending results on Lagrangian chaos and scalar advection, as obtained in [13, 14, 15, 16], to boundary-driven models was discussed in [15, p. 1254].
@@
    Finally, note that a colored-in-time boundary noise–such as replacing the white noise with an Ornstein–Uhlenbeck process–can simplify the check of the integrability condition (1.6). However, in many situations, the white-in-time noise often leads to important insights for the colored-in-time and space case; see [23, Subsection 2.12] and [14, p. 249].

### v2

@@ Abstract.
We show that the Lagrangian flow associated with the stochastic 3D primitive equations (PEs) with nondegenerate noise is chaotic, i.e., the corresponding top Lyapunov exponent is strictly positive almost surely. This result builds on the landmark work by Bedrossian, Blumenthal, and Punshon-Smith [16] on Lagrangian chaos in stochastic fluid mechanics. Our primary contribution is establishing an instance where Lagrangian chaos can be proven for a fluid flow with supercritical energy, a key characteristic of 3D fluid dynamics. For the 3D PEs, establishing the existence of the top Lyapunov exponent is already a challenging task. We address this difficulty by deriving new estimates for the invariant measures of the 3D PEs, which capture the anisotropic smoothing in the dynamics of the PEs. As a by-product of our results, we also obtain the first uniqueness result for invariant measures of stochastic 3D PEs.
@@ 1. Introduction
In this paper, we consider the three-dimensional stochastic primitive equations (PEs) with additive noise on $\mathcal{O}=\mathbb{T}^{2}_{x,y}\times(0,1)$:
@@
Hence, the horizontal velocity field $v_{t}$ is periodic in all directions, while the Dirichlet condition on $w_{t}$ ensures the impermeability of the boundary $\partial\mathcal{O}$. In particular, $v_{t}$ can be considered as a map on the three-dimensional torus $\mathbb{T}^{3}$.
The PEs (1.1) are one of the fundamental models for geophysical flows used to describe oceanic and atmospheric dynamics, where, in a typical region, the vertical scale is much smaller than the horizontal ones. In such domains, the PEs well approximate the 3D Navier-Stokes equations. In the deterministic case (i.e., $Q=0$), rigorous justifications by means of the small aspect ratio limit can be found in [12, 32, 56]. Detailed information on the geophysical background for the various versions of the deterministic PEs can be found in, e.g., [52, 63, 70].
The PEs have been widely studied in recent years, therefore, it is impossible to give a complete overview here. However, we highlight some key references. In the deterministic setting, we point to the seminal work by Lions, Temam, Wang [58, 59], as well as the breakthrough result by Cao and Titi [22] on global well-posedness in $H^{1}$. Extensions in various directions can be found in [34, 35, 43, 44, 49, 53]. In the stochastic setting, global well-posedness was first obtained by Debussche, Glatt-Holtz, Temam and Ziane in [27] (see also [26]). For further developments, the reader is referred to [2, 4, 5, 21, 28, 45] and the references therein.
In contrast to well-posedness, the dynamical properties of solutions to PEs are far less understood. The primary goal of this manuscript is to prove that the Lagrangian trajectories associated with the solution $u_{t}$ of the stochastic PEs (1.1) are chaotic in the sense that they are highly sensitive to initial conditions. This phenomenon, often referred to as Lagrangian chaos, is established in the following.
@@
(1.3) $\frac{\mathrm{d}}{\mathrm{d}t}\phi^{t}(\mathbf{x})=u_{t}(\phi^{t}(\mathbf{x})),\qquad\phi^{0}(\mathbf{x})=\mathbf{x}\in\mathcal{O},$
@@
(1.4) $\lim_{t\to\infty}\frac{1}{t}\log|\nabla\phi^{t}(\mathbf{x})|=\lambda_{+}>0\text{ for all $\mathbf{x}\in\mathcal{O}$ with probability }1.$
The rigorous formulation of the above is given in Theorem 3.5. Equation (1.4) shows that nearby particles are exponentially separated by the flow $\phi^{t}$. The exponential expansion rate $\lambda_{+}$ is usually called the (top) Lyapunov exponent of $\phi^{t}$. Consequences on scalar turbulence of Theorem 1.1 or, more precisely, a version of (1.4) with $\nabla\phi^{t}(\mathbf{x})$ replaced by $(\nabla\phi^{t}(\mathbf{x}))^{-\top}$, are given in Corollary 3.7.
@@
A key element in the proof of Theorem 1.1 is the existence of a unique invariant measure for the stochastic PEs (1.1). Invariant measures play a crucial role in understanding the long-time behaviour of solutions to a given SPDE [25], and their uniqueness has significant implications, including ergodicity and the complete characterisation of long-time dynamics through the invariant measure. For definitions, the reader is referred to Subsection 3.1. While the existence of ergodic measures for PEs was established by Glatt-Holtz, Kukavica, Vicol and Ziane in [37], their uniqueness remains an open problem (see the comments below Theorem C in [19] and the recent work [57]). The main obstacle in establishing uniqueness is the weak bounds available for PEs in their natural state space $H^{\sigma}$ with $\sigma\geq 1$, cf., [37, Section II] and [5, Theorem 3.6 and Remark 3.10]. Here, the weakness is with respect to integrability in the probability space and invariant measures.
In this manuscript, by leveraging on new high-order energy estimates for (1.1) established in Section 5, we are able to overcome these challenges. The following is an informal version and a special case of Proposition 3.2 and Theorem 3.3.
@@
To the best of our knowledge, Theorem 1.2 presents the first uniqueness result for invariant measures of PEs (1.1). Our proof of uniqueness follows the well-established strategy involving strong Feller and irreducibility properties of transition Markov kernels, as used in [31] for the 2D Navier-Stokes equations. For the relevant terminology, we refer the reader to Subsection 3.1 or [25]. In our setting, the high-order energy estimates established in Section 5 play a central role. It is worth noting that, due to the use of the strong Feller approach, we require that the noise acts on all Fourier modes, cf., (3.3) and Subsection 2.2.2. Establishing the uniqueness of invariant measures in the case of degenerate noise–where some Fourier modes are not randomly forced–appears to be extremely challenging due to the above-mentioned weak energy bounds. In particular, known methods, such as asymptotic couplings [38, 42] (see [38, Subsection 3.2] for the two-dimensional variant of (1.1)), asymptotic strong Feller [40], or the $e$-property [51], do not seem to be applicable in this context.
@@ 1.1. Background and contribution
Recall that energy supercriticality is typical of three-dimensional models and produces weak $\Omega$-bounds for the solution of the fluid flow in the well-posedness space. Let us emphasise that the latter also appears in two-dimensional models with boundary noise (see Subsection 1.2 for more details).
The stochastic PEs (1.1) serve as a prototypical example of globally well-posed 3D fluid models in which both of the aforementioned conditions are simultaneously satisfied. While the boundary conditions are explicitly given in (1.2), the energy supercriticality can be seen from a scaling analysis; see [2, Subsection 1.2] for details. It is important to note that the PEs (1.1) possess the scaling of the 3D Navier-Stokes equations, and thus, they are energy supercritical. The energy supercriticality of the PEs is essentially reflected in the weak bounds mentioned above Theorem 1.2, and therefore related difficulties are already present at the level of the fluid equation. Meanwhile, the presence of the impermeability condition (1.2) forces the flow $\phi^{t}$ in (1.3) to live in the non-compact space $\mathcal{O}$. The latter fact will be particularly relevant when dealing with ergodic measures of the so-called Lagrangian process $(u_{t},\phi^{t}(\mathbf{x}))_{t}$, as compactness methods (e.g., Krein-Milman’s theorem) cannot be applied to ensure ergodicity of the invariant measure $\mu\otimes\mathrm{Leb}$ (here, $\mu$ is as in Theorem 1.2). To address this difficulty, we establish auxiliary results on Markov semigroups in Appendix A.
The essential ingredients in the application of the general framework in [16, Section 4] to (1.1) can be (roughly) summarised as follows:
@@
    An integrability condition for the ergodic measures $\mu$ of the fluid model.
@@
    Ergodicity of $\mu\otimes\mathrm{Leb}$ for the Lagrangian process $(u_{t},\phi^{t}(\mathbf{x}))_{t}$.
  4. (4)
    Approximate controllability results for:
    • The Jacobian process $\big{(}u_{t},\phi^{t}(\mathbf{x}),\nabla\phi^{t}(\mathbf{x})\big{)}_{t}.$
    • The projective process $\big{(}u_{t},\phi^{t}(\mathbf{x}),\nabla\phi^{t}(\mathbf{x})\xi/|\nabla\phi^{t}(\mathbf{x})\xi|\big{)}_{t}.$
In the above list, $\mu$ is the unique invariant measure for the stochastic PEs (1.1) provided by Theorem 1.2, $\mathbf{x}\in\mathcal{O}$ and $\xi\in{\mathbb{R}}^{3}\setminus\{0\}$.
@@
(1.7) $w_{t}(\mathbf{x})=-\int_{0}^{z}\nabla_{x,y}\cdot v_{t}(x,y,z^{\prime})\,\mathrm{d}z^{\prime},\ \quad\mathbf{x}=(x,y,z)\in\mathcal{O}.$
Here, $\nabla_{x,y}\cdot$ is the divergence operator with respect to the horizontal variables $(x,y)$; see below for the notation. Consequently, (1.6) requires the integrability of quantities involving second-order derivatives of the horizontal velocity $v_{t}$. However, in the existing literature, only logarithmic bounds on second-order derivatives are available; see [37, Theorem 1.9]. The latter estimates are too weak to ensure (1.6), therefore ensuring the existence of the Lyapunov exponent (1.5).
In this work, we address these issues by using $L^{p}$-estimates and exploiting the anisotropic smoothing effect associated with the nonlinearity $w\partial_{z}v$, where $w$ is as in (1.7). For further details, the reader is referred to Theorem 4.2 and the comments below it. Additionally, we obtain (1.6) with spatial integrability $L^{1}$ replaced by $L^{p}$ for any $p<\frac{4}{3}$. We also expect that the $L^{2}$-case of (1.6) does not hold.
@@
Before discussing the subsequent item, let us comment on other possible choices of smoothing for the Markov kernel of the Lagrangian process. In the recent work [23], the authors proved that in the context of the 2D Navier-Stokes equations the results in [16] (as well as subsequent works [13, 14, 15]) can be extended to the case where the noise acts on only finitely many Fourier modes (i.e., degenerate noise). This extension is achieved through the use of the asymptotic strong Feller property, which was introduced in [40] and used to show unique ergodicity of the 2D Navier-Stokes equations with degenerate noise. However, as noted below Theorem 1.2, it remains unclear how to establish asymptotic strong Feller for the stochastic PEs (1.1) itself in the case of degenerate noise. Therefore, the extension presented in [23] of the results in [16] to the case of degenerate noise does not currently seem applicable in the context of the stochastic PEs (1.1).
(3): Although the invariance of the measure $\mu\otimes\mathrm{Leb}$ under the Lagrangian dynamics $(u_{t},\phi^{t}(\mathbf{x}))_{t}$ follows directly from the divergence-free condition on $u_{t}$, the question of ergodicity is more delicate. Unlike in [16], due to the non-compactness of $\mathcal{O}$, it is difficult to obtain uniqueness of invariant measures for the Lagrangian process $(u_{t},\phi^{t}(\mathbf{x}))_{t}$, and hence, ergodicity of $\mu\otimes\mathrm{Leb}$. Indeed, in [16], the compactness of the underlined domain allows one to invoke the Krein-Milman theorem along with the uniqueness and extremality of ergodic measures (see, e.g., [25, Proposition 3.2.7]). Moreover, this issue cannot be solved by considering the flow $\phi^{t}$ on the compact set $\overline{\mathcal{O}}$. Indeed, in the latter situation, it is possible to construct ergodic measures for the Lagrangian process that are supported on $\mathrm{supp}\,\mu\times\partial\mathcal{O}$. Combined with the Krein-Milman theorem and the invariance of $\mu\otimes\mathrm{Leb}$, this implies the existence of multiple ergodic measures for the Lagrangian process $(u_{t},\phi^{t}(\mathbf{x}))_{t}$ where $\phi^{t}$ is considered on $\overline{\mathcal{O}}$. Nevertheless, by employing a topological argument along with the strong Feller property of the process $(u_{t},\phi^{t})_{t}$ with $\mathbf{x}\in\mathcal{O}$, we can still establish ergodicity of $\mu\otimes\mathrm{Leb}$ via the results in Appendix A.
@@
To conclude, let us emphasise that, as in [16], we will obtain a stronger result than (1.4) in Theorem 1.1. Specifically, using the random multiplicative ergodic theorem [50, Theorem III.1.2] (also used in [16], see Proposition 3.16 therein), one can derive from Theorem 1.1 the following stronger result: For all $\xi\in{\mathbb{R}}^{3}\setminus\{0\}$,
(1.8) $\lim_{t\to\infty}\frac{1}{t}\log|\nabla\phi^{t}(\mathbf{x})\xi|=\lambda_{+}>0\text{ for all $\mathbf{x}\in\mathcal{O}$ with probability }1.$
@@
    Uniqueness of invariant measures for the projective process, whose projection onto the first two components coincides with $\mu\otimes\mathrm{Leb}$.
As in (3), we are unable to prove uniqueness of the invariant measure for the projective process in general, but only within the class of measures whose projection onto the first two components coincides with $\mu\otimes\mathrm{Leb}$. This limitation again stems from the non-compactness of $\mathcal{O}$. The claim (5) will be established by proving uniqueness of ergodic measures for the projective process, together with the ergodicity of $\mu\otimes\mathrm{Leb}$ for the Lagrangian dynamics. The uniqueness of ergodic measures is obtained via the strong Feller property of the projective process, as shown in Propositions 3.11 and 3.12. We emphasise that the strong Feller property established here is a stronger result than what is required in (2) from the previous list.
@@ 1.2. Future challenges
  • Mixing and enhanced dissipation of passive scalars. As shown in the works [13, 14], the positivity of the top Lyapunov exponent of the flow map associated with a stochastic fluid model often implies mixing and enhanced dissipation results for passive and diffusive scalars, respectively. It is important to note that these results are stronger than the exponential growth for passive scalars presented in Corollary 3.7. While the positivity of the top Lyapunov exponent for the induced flow map is a key element in [13, 14], these results also rely on the existence of Lyapunov functions that provide quantitative control over high deviations of the velocity $u_{t}$, as discussed in [14, Subsection 2.4]. The construction of such Lyapunov functions remains an open problem for stochastic PEs (1.1).
  • Other geophysical flows: Boundary noise. The interaction between the wind and the ocean surface plays a crucial role in ocean dynamics, see [36, 63, 64]. In the case of highly oscillating wind, and under the assumption of a time-scale separation, the wind’s effect can be idealised as noise acting on the boundary. Recent well-posedness results for fluid dynamics models with boundary noise are presented in [3, 6, 18]. The physical relevance of extending results on Lagrangian chaos and scalar advection, as obtained in [13, 14, 15, 16], to boundary-driven models was discussed in [15, p. 1254].
@@
    Finally, note that a colored-in-time boundary noise–such as replacing the white noise with an Ornstein-Uhlenbeck process–can simplify the proof of the condition (1.6). However, in many situations, the white-in-time noise often leads to important insights for the colored-in-time and space case; see [23, Subsection 2.12] and [14, p. 249].